# ECAMP: Entity-centered Context-aware Medical Vision Language Pre-training

### v3

@@ Abstract
Despite significant advancements in medical vision-language pre-training, existing methods have largely overlooked the inherent entity-specific context within radiology reports and the complex cross-modality contextual relationships between text and images. To close this gap, we propose a novel Entity-centered Context-aware Medical Vision-language Pre-training (ECAMP) framework, which is designed to enable a more entity-centered and context-sensitive interpretation of medical data. Utilizing the recent powerful large language model, we distill entity-centered context from medical reports, which enables ECAMP to gain more effective supervision from the text modality. By further pre-training our model with carefully designed entity-aware, context-enhanced masked language modeling and context-guided super-resolution tasks, ECAMP significantly refines the interplay between text and image modalities, leading to an enhanced ability to extract entity-centered contextual features. Besides, our proposed multi-scale context fusion design also improves the semantic integration of both coarse and fine-level image representations, prompting better performance for multi-scale downstream applications. Combining these components leads to significant performance leaps over current state-of-the-art methods and establishes a new standard for cross-modality learning in medical imaging, whose effectiveness is demonstrated by our extensive experiments on various tasks including classification, segmentation, and detection across several public datasets. Code and models are available at https://github.com/ToniChopp/ECAMP.
@@
medical vision-language pre-training masked modeling
@@ 1 Introduction
By virtue of its exceptional accuracy and efficiency, deep learning (DL) is increasingly pivotal in medical image analysis [64, 39, 14, 13, 46]. However, precision is contingent upon large-scale and high-quality annotations, which require an extensive number of experienced medical professionals, rendering it a costly endeavor. Fortunately, the diagnosis and experiential knowledge [40] of radiologists are preserved in the form of paired reports and images, which provide valuable semantic context for training DL-based models. More notably, these supervised signals are relatively accessible at little to no additional costs.
@@
Figure 1: Our ECAMP obtains medical knowledge from ChatGPT and experts respectively, centered on entities including diseases and symptoms, etc. Context distillation asks ChatGPT to distillate precise content for entities, while Context-enhanced MLM predicts the masked descriptors close to entities and addresses the imbalance difficulty caused by doctors’ habits. Further, ECAMP develops Multi-scale Context Fusion to learn global and local projections simultaneously for multi-granularity downstream tasks.
@@
Cutting-edge algorithms [1] in medical vision-language pre-training (Med-VLP) primarily leverage the information-rich medical report-image pairs to learn generic representations from two aspects: contrastive-based methods align the text and image features at both global [61] and local [17, 49] levels; while reconstruction-based methods generate the masked words [11] and image patches [15] using cross-modality information [63]. Nevertheless, these methods still exhibit limitations in harnessing the potential of supervision in the report context. First, the linguistic challenge [22] posed by complex biomedical context [29] impeds the effectiveness of pre-training [2]. Second, radiologists often include diagnoses for common diseases (e.g., pneumothorax) in their reports, even when these diseases are negative [10]. This results in a significant context imbalance issue [2]: the majority of disease diagnoses are negative [^1] Statistical results of the imbalanced diagnosis descriptors can be found in the supplementary materials., introducing an adverse bias into the optimization of masked language modeling (MLM) [11].
In this paper, we introduce Entity-Centered Context-Aware Medical vision-language Pre-training (ECAMP), a robust pre-training framework for radiology. As shown in Fig. 1, ECAMP consists of four simple-yet-effective components: 1) entity-aware context distillation, 2) entity-centered context-enhanced MLM, 3) context-guided super-resolution, and 4) multi-scale context fusion. Superior to PRIOR [7] which substitutes complex biomedical sentences with prototypes, we propose entity-aware context distillation to distill context from Large Language Model in the form of text, leveraging its strong power in linguistic comprehensive and summarizing [36]. Specifically, we use ChatGPT [3] to summarize the complex reports into succinct expressions with entity diagnosis (like “There may be pneumonia.”), which facilitates more effective supervision of the complex reports in pre-training. Furthermore, based on our consultations with medical professionals [10], radiologists typically express diagnostic findings by using descriptors close to Medical Subject Headings (MeSH) [30] entities (e.g., pneumonia). Accordingly, we propose entity-centered context-enhanced MLM, employing a fixed masking strategy to optimize the pre-trained model to predict these key descriptors. To tackle the imbalance problem, we further add an effective re-balanced factor for the minority of the positive descriptors, which prevents the model from ignoring the critical positive diagnoses in reports.
@@
Moreover, super-resolution exhibits promise as a self-supervised proxy task [63]. Given that critical pathology usually occupies small patches, we hypothesize that learning to reconstruct these subtle cues helps to better understand pathology and hence leads to better performance. Fortuitously, leveraging the concise distilled reports, we can retrain patch-level contrastive-based methods [17] to procure precise entity-specific attention maps. These maps facilitate our proposed context-guided super-resolution to learn the high-resolution characteristics of the target pathology with broader contextual information, thereby enabling a more holistic and accurate interpretation of medical images.
@@
In addition, we find current cross-modal reconstruction-based methods typically employ either global average pooling (GAP) [63] or cross-attention [6] to integrate global and local vision representations to assist MLM. In contrast, we introduce multi-scale context fusion to aggregate global and local projection of pre-trained vision encoder simultaneously. Our motivation is to learn both global and local informative representations, which can be generalized to a wide range of multi-scale downstream tasks, including classification, segmentation, and detection.
@@
We pre-train ECAMP on MIMIC-CXR [19] dataset consisting of image-report pairs. Extensive experiments are conducted on multi-scale downstream tasks (including classification, segmentation, and detection) across multiple public datasets, e.g., ChestX-ray14 [51], CheXpert [18], RSNA Pneumonia [43], COVIDx [50] and SIIM Pneumothorax [44] to demonstrate the superiority of ECAMP compared with a plethora of competitive state-of-the-art Med-VLP methods with a significant performance gap, whose readily accessible components prove valuable in fortifying the establishment of a foundational medical model.
@@ 2.1 General Vision Language Pre-training
Figure 2: The overall framework of our ECAMP. Initially, entity-aware context distillation succinctly encapsulates precise diagnosis for each entity. Then entity-centered context-enhanced MLM masks the two descriptive words (underlined) preceding the entity, employing a re-balanced vector to direct the model’s attention towards critical minority-positive expressions. To further enhance the learning of pathology vision contexts, context-guided super-resolution utilizes an entity-specific attention map. Finally, we introduce a novel multi-scale context fusion block (highlighted in cyan) to concurrently capture representative global and local features.
@@ 2.2 Medical Vision Language Pre-training
Recent Med-VLP methods can be categorized into two distinct types: report-supervised cross-modal alignment pre-training and reconstruction-based self-supervised pre-training. The groundbreaking works of the former type such as ConVIRT [61], REFERS [62], and CheXZero [47] pre-train the model by directly maximizing mutual information between the global representations. GLoRIA [17], SAT [31], and MGCA [49] propose to align fine-grained features of paired image patches and words. PRIOR [7] and BioVIL [2] attempt to comprehend intricate medical reports. MedKLIP [54] and KAD [60] extract the medical-related information using a triplet extraction module as additional supervised signals. Inspired by BERT [11] and MAE [15], the latter type learns representation by cross-modal context reconstruction task. M3AE [6] utilizes cross-attention for integrating multi-modal information to reconstruct masked tokens. While MRM [63] directly applies global average pooling for fusing visual features for masked text modeling. Further, MPMA [57] introduces a memory-augmented cross-modal fusion module to fully integrate visual information to assist report reconstruction.
Advanced efforts try to combine contrastive-based and reconstruction-based pre-training. CMITM [5] contrastively aligns global report and image representations after pre-training with MRM [63], while MedIM [56] boosts MGCA [49] with attention guided masked image modeling. Notwithstanding the enhancements, the aforementioned approaches are constrained by the intricacies of comprehending biomedical reports and the imbalance in diagnoses, which has spurred our pursuit of developing ECAMP. Med-Unic [48] and Med-MLLM [32] further improve the performance by leveraging more pre-training datasets.
@@ 3 Method
In this section, we first introduce how ECAMP leverages context distillation to distill distinct context from LLM in Sec. 3.1. Then, we illustrate the basic masked image modeling (MIM) [15] and the proposed entity-centered context-enhanced MLM in Sec. 3.2. Next, Sec. 3.3 shows the process of context-guided super-resolution. Finally, we clarify the detailed approaches of multi-scale context fusion in Sec. 3.4 to integrate both global and local vision features for MLM.
@@ 3.1 Context Distillation
Given an image $I^{o}\in\mathbb{R}^{H^{o}\times W^{o}}$ and its original paired report $T_{o}$, Our target is to extract concise diagnosis for diseases from $T_{o}$. Accordingly, we select the set of entities $R=\{r^{1},r^{2},\ldots,r^{k}\}$ from MeSH [30], representing the diseases of interest in chest radiography. For each original report $T_{o}$, we identify entities $\mathbf{r}_{o}$ mentioned in the report through string matching. Next, prompt $p$ is designed to require LLM to comprehend the complex biomedical context and summarize the diagnosis for disease entities $\mathbf{r}_{o}$. Furthermore, to make the summarization of LLM stable, we leverage in-context learning [34] to construct an one-shot [53] sample $p^{\prime}$ in the final prompt [^2] The selected entities of disease and detailed prompt can be found in the Appendix.B and Appendix.C, respectively. $p=P(T_{o},\mathbf{r}_{o},p^{\prime})$. The distilled report $T_{d}$ is finally computed as follows:
@@ 3.2 Masked Image and Language Modeling
Masked image modeling We first down-sample the original image $I^{o}$ into high-resolution $I^{h}\in\mathbb{R}^{H^{h}\times W^{h}}$ and low-resolution $I^{l}\in\mathbb{R}^{H^{l}\times W^{l}}$ input. The low-resolution image $I^{l}$ is split into $N_{p}$ patches with size $P\times P$. Following MAE [15], we mask a large ratio ($75\%$) of the $N_{p}^{I}$ low-resolution patches, resulting in the unmasked patches $I_{u}=\{{I}_{u}^{1},I_{u}^{2},\ldots,I_{u}^{N_{u}^{I}}\}$ and masked patches $I_{m}=\{{I}_{m}^{1},I_{m}^{2},\ldots,I_{m}^{N_{m}^{I}}\}$.Then the patch embedding $E_{p}$ is obtained by linearly projecting the flattened unmasked patches $I_{u}$. Vision transformer [12] (ViT) is chosen as the vision encoder $E_{I}$, which projects the patch embedding $E_{p}$ along with its positional embedding $E_{pos}$ as the vision feature $f_{v}=E_{I}(I_{u},E_{pos})$. In the image decoding procedure, $f_{v}$ is fed into an image decoder $D_{I}$ after recovering to the original $N_{p}$ size, resulting in reconstructed image $\hat{I^{l}}=D_{I}(E_{I}(I_{u}))$. We choose mean square error (MSE) loss to optimize the reconstructed image:
@@
A high text masking ratio challenges the model to reconstruct the reports by encoding the vision context thoroughly and accurately. Accordingly, a novel multi-scale context fusion module (MSCF, see Sec. 3.4) is proposed to integrate the vision feature and text embedding $f_{f}=\mathrm{MSCF}(f_{v},E_{t})$. The reconstructed text $\hat{T}$ is generated by decoding the fused feature $f_{f}$ using a BERT-based [11] text decoder $\hat{T}=D_{t}(f_{f})$. The MLM loss for each masked word $t\in\left[T_{m},T_{a}\right]$ is defined as $\mathcal{L}^{t}_{\mathrm{MLM}}(\hat{t},t)=-\mathrm{log}(P(\hat{t}=t))$.
@@
Moreover, we simply categorize $T_{a}$ into two groups: negative descriptions $T_{neg}$ and others $T_{oth}$, with $N^{T}_{neg}$ and $N^{T}_{oth}$ words, respectively. Please note that since the expressions of positive descriptors are diverse, we divide them into others $T_{oth}$. To focus on modeling the minority-positive disease descriptions instead of majority-negative words, we multiply the loss of $T_{neg}$ by a small number $\lambda_{neg}$, and the loss of $N^{T}_{oth}$ by $\lambda_{oth}$, where
@@
According to the statistical results in Appendix.A, the most negative descriptor “there is no" occurs approximately 20×20$\times$20 × more often than the most positive descriptor “there may be.", we set $\lambda_{neg}=0.05$. Please note that $\lambda_{oth}$ is important to make sure that the model pays an enough attention to accurately reconstruct the critical “descriptive words” $T_{a}$. The final loss of entity-centered context-aware MLM is defined:
@@
Method CXR14 (AUC) CXP (AUC) RSNA (AUC) COVIDx (ACC)

### v4

@@ Abstract
Despite significant advancements in medical vision-language pre-training, existing methods have largely overlooked the inherent linguistic complexity and imbalanced isssue within medical reports, as well as the complex cross-modality contextual relationships between texts and images. To close this gap, we propose a novel Entity-centered Context-aware Medical Vision-language Pre-training (ECAMP) framework, which establishes a more entity-centered, context-sensitive, and balanced understanding of medical reports to effectively pre-train the vision encoder. We first distill entity-centered context from medical reports utilizing large language models, enabling ECAMP to draw more precise supervision from the text modality. By further incorporating entity-aware re-balanced factor and descriptor masking strategies into masked languange modeling, ECAMP significantly enhances the knowledge of entities within the reports. A context-guided super-resolution task is proposed alongside a multi-scale context fusion design to improve the semantic integration of both coarse and fine-level image representations, which prompts better performance for multi-scale downstream applications. ECAMP integrates these innovations together, leading to significant performance leaps over current state-of-the-art methods and establish a new standard for cross-modality pre-training in medical imaging. The effectiveness of ECAMP is demonstrated by extensive experiments on various domains and organs, which achieves cutting-edge results on multiple tasks including classification, segmentation, and detection across 5 public chest X-ray and 4 fundoscopy datasets respectively.
@@
[…]
journal: Medical Image Analysis
@@ 1 Introduction
By virtue of its exceptional accuracy and efficiency, deep learning (DL) is increasingly pivotal in medical image analysis [74, 44, 14, 13, 53]. However, precision is contingent upon large-scale and high-quality annotations, which require an extensive number of experienced medical professionals, rendering it a costly endeavor. Fortunately, the diagnosis and experiential knowledge [45] of doctors are preserved in the form of paired reports and images, which provide valuable semantic context for training DL-based models. More notably, these supervised signals are relatively accessible at little to no additional costs.
@@
Fig. 1: Our ECAMP is designed to learn representative and multi-scale features from complex and imbalanced medical reports, extracting medical knowledge from both ChatGPT and experts, with a particular focus on entities such as diseases, symptoms, and others. Context distillation prompts ChatGPT to distillate precise content for entities, thereby reducing the complexity of reports, while Context-enhanced MLM predicts the masked descriptors close to entities and addresses the imbalance difficulty caused by doctors’ habits. Context-guided Super Resolution focuses the model on anatomy area, concentrating the target pathology textures and locations. Further, ECAMP develops Multi-scale Context Fusion to learn global and local projections simultaneously to support multi-granularity downstream tasks.
@@
Cutting-edge algorithms [1, 66] in medical vision-language pre-training (Med-VLP) primarily leverage the information-rich medical report-image pairs to learn generic representations from two aspects: contrastive-based methods align the text and image features at both global [71] and local [17, 56, 34] levels; while reconstruction-based methods generate the masked words [11] and image patches [15] using cross-modality information [73].
Nevertheless, these methods still exhibit limitations in harnessing the potential of supervision in the context of the report. First, the linguistic challenge [25] posed by complex biomedical context [33] impedes the effectiveness of pre-training [2]. Second, doctors often include diagnoses for common diseases (e.g., pneumothorax) in their reports, even when these diseases are negative [10]. This practice leads to a significant context imbalance issue [2]: the majority of disease diagnoses are negative. We extract and count the two adjective words before disease entities (defined as “descriptors” ) and plot the statistics of these descriptors in Fig. 2 (b)., which clearly indicate an imbalance: the ’is no’ descriptor is dominant and the imbalance ratio between negative (e.g., “is no”) and positive (e.g., “there is”) descriptors reaches 20:1. This imbalance injects an adverse bias into the optimization of masked language modeling (MLM) [11] and complicates the model’s ability to accurately learn the semantics of the positive diseases.
In this paper, we introduce Entity-Centered Context-Aware Medical vision-language Pre-training (ECAMP), a robust pre-training framework for 2D medical images. As shown in Fig. 1, the main contribution of ECAMP is pre-training representative and multi-scale features from complex and imbalanced medical reports through a cohesive integration of four synergistic modules. ECAMP consists of four simple-yet-effective components: 1) entity-aware context distillation, 2) entity-centered context-enhanced MLM, 3) context-guided super-resolution, and 4) multi-scale context fusion. We propose a cohesive integration of four modules, which contribute from different perspectives with close collaboration, emphasizing an entity-centered and context-aware design.
Fig. 2: The statistics of high-frequency entities and descriptors in the original reports of MIMIC-CXR [21]. We merge other low-frequent words as “others”. Our selected entities account for approximately 48.6% of the overall occurrences. The imbalance ratio between negative (e.g., “is no”) and positive (e.g., “there is”) descriptors reaches to 20:1.
Unlike PRIOR [7] that substitutes complex biomedical sentences with prototypes, we propose entity-aware context distillation to distill semantic context using Large Language Model, leveraging its strong power in linguistic comprehensive and summarization [41]. Based on our consultations with medical professionals [10], doctors typically express diagnostic findings by using descriptors close to Medical Subject Headings (MeSH) [35] entities (e.g., pneumonia). Our key observations related to medical reports motivate the need for better distilling the semantic contexts related to clinically-relevant entities from reports. Specifically, we use ChatGPT [3] to summarize the complex reports into succinct expressions with entity diagnosis (like “There may be pneumonia.”). We construct an entity list, including diseases and symptoms, and further generate more precise and simplified descriptions of disease existence and severity, which facilitate more effective supervision of the complex reports in pre-training. To the best of our knowledge, we are the first to leverage the power of large language models for medical vision-language pre-training, which attempts to explicitly address the challenge of information overload in clinically dense reports. The distilled context is not just a simplified report; it contains more precise descriptions of the existence and severity of diseases, enabling our model to learn more precise anatomical locations and disease-related features from the complex reports.
Despite that multi-modal MLM has been investigate in recent Med-VLP methods such as M3AE [6] and MRM [73], they randomly mask portions of the report for construction, which overlook the importance and imbalance of the descriptive words beyond the entity. Accordingly, we propose entity-centered context-enhanced MLM strategy. The descriptor masking employs a fixed masking strategy to optimize the pre-trained model to predict these key descriptors for better modeling the semantics of entities and context. To tackle the imbalance problem (as shown in Fig 2 (b).), we further add an effective re-balancing factor for less-appearing positive descriptors, which prevents the model from ignoring the critical positive diagnoses in reports.
@@
Moreover, traditional MIM methods, such as MAE [15], treat each patch of the image equally during loss computation. However, lesions usually locate at small area of imaging. Super-resolution exhibits promise as a lesion focus and reinforcement of masked image modeling in self-supervised proxy task [73]. Given that critical pathology usually occupies small patches, we hypothesize that learning to reconstruct these subtle cues help to better understand pathology and hence lead to better performance. Fortuitously, leveraging the concise distilled reports, we can retrain patch-level contrastive-based methods [17] to procure precise entity-specific attention maps. These maps facilitate our proposed context-guided super-resolution to learn the high-resolution characteristics of the target pathology with broader contextual information, thereby enabling a more holistic and accurate interpretation of medical images based on the entities and context.
@@
In addition, we find current cross-modal reconstruction-based methods typically employ either global average pooling (GAP) [73] or cross-attention [6] to integrate global and local vision representations to assist MLM. In contrast, we introduce multi-scale context fusion to aggregate global and local projection of pre-trained vision encoder simultaneously. Multi-scale feature aggregation is definitely a trend for Med-VLP, but different methods has special designs which fit their frameworks. GLoRIA [17], which uses contrastive learning to aggregate and align multi-scale features in two different branch. Instead, the only supervision which contains diagnosis information in our framework comes from masked report reconstruction, the gradient is propagated to local and global features via addition and duplication simultaneously, and train the multi-scale features. Our motivation is to learn both global and local informative representations, which can be generalized to a wide range of multi-scale downstream tasks, including classification, segmentation, and detection.
Fig. 3: The overall framework of our ECAMP. Initially, entity-aware context distillation succinctly encapsulates precise diagnosis for each entity. Then entity-centered context-enhanced MLM masks the two descriptive words (underlined) preceding the entities, employing a re-balancing vector to direct the model’s attention towards critical minority-positive expressions. To further enhance the learning of pathology vision contexts, context-guided super-resolution utilizes an entity-specific attention map. Finally, we introduce a novel multi-scale context fusion block (highlighted in cyan) to concurrently capture representative global and local features.
@@
To comprehensively and thoroughly validate the effectiveness, We pre-train our medical pre-training method ECAMP on MIMIC-CXR [21] and FFA-IR [29] dataset consisting of image-report pairs for both Chest X-rays and fundoscopies. Extensive experiments are conducted on multi-scale downstream tasks (including classification, segmentation, and detection) across multiple public datasets. The experimental results across various domains and organs demonstrate the superiority and generalization of ECAMP to outperform a plethora of competitive state-of-the-art (SOTA) Med-VLP methods with significant performance gaps. In sum, ECAMP with its readily accessible components prove valuable in fortifying the establishment of pre-training framework for medical foundation model. Codes and models are available at https://github.com/ToniChopp/ECAMP.
@@ 2.2 Medical Vision Language Pre-training
Recent Med-VLP methods can be categorized into two distinct types: report-supervised cross-modal alignment pre-training and reconstruction-based self-supervised pre-training. The groundbreaking works of the former type such as ConVIRT [71], REFERS [72], and CheXZero [54] pre-train the model by directly maximizing mutual information between the global representations. GLoRIA [17], SAT [36], and MGCA [56] propose to align fine-grained features of paired image patches and words. PRIOR [7] and BioVIL [2] attempt to comprehend intricate medical reports. MedKLIP [61] and KAD [70] extract the medical-related information using a triplet extraction module as additional supervised signals. FLAIR [50] integrates the expert’s domain knowledge in the form of descriptive textual prompts. RETFound [75] provides a basis for better label-effcient model adaptation. Inspired by BERT [11] and MAE [15], the latter type learns representation by cross-modal context reconstruction task. M3AE [6] utilizes cross-attention for integrating multi-modal information to reconstruct masked tokens. While MRM [73] directly applies global average pooling for fusing visual features for masked text modeling. Further, MPMA [67] introduces a memory-augmented cross-modal fusion module to fully integrate visual information to assist report reconstruction.
Advanced efforts try to combine contrastive-based and reconstruction-based pre-training. CMITM [5] contrastively aligns global report and image representations after pre-training with MRM [73], while MedIM [64] boosts MGCA [56] with attention guided masked image modeling. Med-Unic [55] and Med-MLLM [37] further improve the performance by leveraging more pre-training datasets. Pair-Aug [63] proposes pairwise augmentation to improve performance without introducing extra datasets. Notwithstanding the enhancements, the aforementioned approaches are constrained by the intricacies of comprehending biomedical reports and the imbalance in diagnoses, which has spurred our pursuit of developing ECAMP.
@@ 2.3 Large Language Models (LLMs)
Table 1: Our selected entities. We select 44 entities based on their importance and frequency of appearance in radiology reports as comprehensively as possible.
Entities
abnormality abscess aerate aorta
atelectasis bronchiectasis calcification cardiomediastinal
cardiomegaly catheter chf collapse
congestion consolidation contour COPD
deformity dilation distention edema
effusion embolism emphysema engorgement
fibrosis fracture granuloma hernia
hilar hyperinflate hemidiaphragm infiltrate
mass nodule obscure opacity
perihilar pneumonia pneumothorax sarcoidosis
silhouette thickening tuberculosis vasculature
@@ 3 Method
In this section, we take CXRs and corresponding reports as an example to illustrate our proposed method. We first introduce how ECAMP leverages context distillation to distill distinct context from LLM in section 3.1. Then, we illustrate the basic masked image modeling (MIM) [15] and the proposed entity-centered context-enhanced MLM in section 3.2. Next, section 3.3 shows the process of context-guided super-resolution. Finally, we clarify the detailed approaches of multi-scale context fusion in section 3.4 to integrate both global and local vision features for MLM.
@@ 3.1 Context Distillation
Given an image $I^{o}\in\mathbb{R}^{H^{o}\times W^{o}}$ and its original paired report $T_{o}$, our target is to extract concise diagnosis for diseases from $T_{o}$. Accordingly, we select the set of entities $R=\{r^{1},r^{2},\ldots,r^{k}\}$ from MeSH [35], representing the diseases of interest in chest radiography. For each original report $T_{o}$, we identify entities $\mathbf{r}_{o}$ mentioned in the report through string matching. We manually curate a list of 44 entities, as shown in Table 1, including diseases, symptoms and anatomical locations. These entities are sourced from MeSH [35] and selected based on their importance and the frequency of appearance in radiology reports. As illustrated in Fig. 2 (a)., the 44 entities selected for this study account for approximately 48.6% of the overall occurrences. Next, prompt $p$ is designed to require LLM to comprehend the complex biomedical context and summarize the diagnosis for disease entities $\mathbf{r}_{o}$. Furthermore, to make the summarization of LLM stable, we leverage in-context learning [39] to construct an one-shot [60] sample $p^{\prime}$ in the final prompt $p=P(T_{o},\mathbf{r}_{o},p^{\prime})$. Table 2 displays the prompt utilized for report distillation. By inputting the radiology report as the content to ChatGPT [41] with this specific prompt, we obtain a distilled report that includes more detailed semantics regarding the existence and severity of entities. Specifically, by using GPT3.5-turbo API [41] and setting temperature with 0 to generate 10,000 distilled reports, to save API call, we train a Vicuna7B [8] to further distill all the reports. Examples of these distilled reports are presented in Table 3.  [^3] A medical student is tasked with evaluating the accuracy of the context distillation. The average distillation accuracy for ChatGPT 3.5 turbo is 93.6%, while the accuracy for our Vicuna 7B is 93.2%. The distilled report $T_{d}$ is finally computed as follows:
@@ 3.2 Masked Image and Language Modeling
Masked image modeling We first decompose the original image $I^{o}$ into high-resolution $I^{h}\in\mathbb{R}^{H^{h}\times W^{h}}$ and low-resolution $I^{l}\in\mathbb{R}^{H^{l}\times W^{l}}$ input. The low-resolution image $I^{l}$ is split into $N_{p}^{l}$ patches with size $P\times P$. Following MAE [15], we mask a large ratio ($75\%$) of the $N_{p}^{I}$ low-resolution patches, resulting in the unmasked patches $I_{u}=\{{I}_{u}^{1},I_{u}^{2},\ldots,I_{u}^{N_{u}^{I}}\}$ and masked patches $I_{m}=\{{I}_{m}^{1},I_{m}^{2},\ldots,I_{m}^{N_{m}^{I}}\}$.Then the patch embedding $E_{p}$ is obtained by linearly projecting the flattened unmasked patches $I_{u}$. Vision transformer [12] (ViT) is chosen as the vision encoder $E_{I}$, which projects the patch embedding $E_{p}$ along with its positional embedding $E_{pos}$ as the vision feature $f_{v}=E_{I}(I_{u},E_{pos})$. In the image decoding procedure, $f_{v}$ is fed into an image decoder $D_{I}$ after recovering to the original $N_{p}$ size, resulting in reconstructed image $\hat{I^{l}}=D_{I}(E_{I}(I_{u}))$. We choose mean square error (MSE) loss to optimize the reconstructed image:
@@
Table 2: Our prompt used for report distillation is designed to include both an indication and an example. Providing ChatGPT [41] with an example of the task helps enhance its performance in generating distilled reports.
Messages = [ { "role": "system", "content": f""" You are a knowledgeable and veteran doctor. """ }, { "role": "user", "content": f""" Please help me analysis the medical reports and conclude them briefly. Now, I will give you a medical report, as well as the mentioned entities. Please write brief and clear conclusions according to the reports as the following format: ’There is [a] [b].’ when you are sure whether the entity exists in the report, or ’There may [a] [b].’, when you are not sure whether the entity exists in the report. [a] represents adjective words describing the severeness or existence of the entity [b]. Please generate conclusions for the entities mentioned above one by one, according to the format and do not generate other words. Please keep only one entity in a sentence, there is no need of using ’and’ or ’or’ to connect two or more words. """ }, { "role": "user", "content": f"""
—Example—-
Report: As compared to _, the lung volumes have slightly decreased. Signs of mild over inflation and moderate pleural effusion persist. Elongation of the descending aorta. Entities: aorta, inflation, effusion Conclusion: There is moderate pleural effusion. There is mild over inflation. There is descending aorta. —Example END—- Given the report: {report}
Entities: {entity}.
Conclusion: """ },]
Fig. 4: The framework of multi-scale context fusion. The addition operation (highlighted in cyan) concurrently fuses the text feature, local aggregated vision feature, and global vision feature. The red arrows indicate the propagation of gradients, which simultaneously flow toward both global and local vision features.
@@
Table 3: Examples of report distillation are presented. For each report, we showcase how the semantics of our selected entities are distilled by ChatGPT and our Vicuna-7B. The entities are highlighted in multiple colors for clarity. The distilled reports provide more specificity regarding the existence of entities.
Original There is mild left base atelectasis seen on the frontal view without clear correlation on the lateral view. No definite focal consolidation is seen. There is no pleural effusion or pneumothorax. The aorta is slightly tortuous. The cardiac silhouette is not enlarged. There is no overt pulmonary edema, mild left base atelectasis. otherwise, no acute cardio pulmonary process.
ChatGPT There is mild left base atelectasis. There is a slightly tortuous aorta. There is no pleural effusion. There is no pneumothorax. There is no overt pulmonary edema. There is no acute cardio pulmonary process.
Vicuna-7B There is mild left base atelectasis. There is no pleural effusion. There is no pneumothorax. There is no definite focal consolidation. There is no enlarged cardiac silhouette. There is no overt pulmonary edema.
Original In comparison with the study of _, the increased opacification at the right base has essentially cleared with better inspiration. Cardiac silhouette remains at the upper limits of normal in size and there is again tortuosity of the aorta without vascular congestion or pleural effusion. Bi-apical changes especially on the right are stable.
ChatGPT There is tortuosity of the aorta. There is no pleural effusion. There is no vascular congestion. There is no change in the bi-apical changes. There is stable increased opacification at the right base. The cardiac silhouette is at the upper limits of normal in size.
Vicuna-7B There is tortuosity of the aorta. There is no pleural effusion. There is no vascular congestion.
Original Pa and lateral chest radiographs again demonstrate severe hyperinflation and diffuse bronchiectasis. There is no focal consolidation, pleural effusion or pneumothorax. The cardiomediastinal silhouette is stable. No acute cardio pulmonary process. Findings were relayed to Dr _.
ChatGPT There is severe hyperinflation. There is diffuse bronchiectasis. There is no focal consolidation. There is no pleural effusion. There is no pneumothorax. The cardiomediastinal silhouette is stable. There is no acute cardio pulmonary process.
Vicuna-7B There is severe hyperinflation. There is diffuse bronchiectasis. There is no focal consolidation. There is no pleural effusion. There is no pneumothorax.
Original The lung volumes are normal. Mild cardiomegaly which is stable. Normal hilar and mediastinal structures. No pneumonia. No pulmonary edema . No pleural effusion. Status post cabg with aligned median stereotomy wires and normal location of surgical clips. Status post right lung surgery with surgical material seen. Mild cardiomegaly. No evidence of pneumonia.
ChatGPT There is normal hilar. There is mild cardiomegaly. There is no pneumonia. There is no pulmonary edema. There is no pleural effusion. There is status post CABG with aligned median sternotomy wires. There is status post right lung surgery with surgical material. The lung volumes are normal.
Vicuna-7B There is normal hilar. There is mild cardiomegaly. There is no pneumonia. There is no pulmonary edema. There is no pleural effusion.
Original 3 radiographs are provided. On the film, 3 the tip of the dobbhoff catheter projects over the proximal parts of the stomach. The course of the device is unremarkable. No complications. Unchanged appearance of the lung parenchyma the pleural and the heart.
ChatGPT There is no complication with the Dobbhoff catheter. There is unchanged appearance of the lung parenchyma. There is no pleural abnormality. There is no heart abnormality. There is no disease.
Vicuna-7B There is no disease.
@@
A high text masking ratio challenges the model to reconstruct the reports by encoding the vision context thoroughly and accurately. Accordingly, a novel multi-scale context fusion module (MSCF, see section 3.4) is proposed to integrate the vision feature and text embedding $f_{f}=\mathrm{MSCF}(f_{v},E_{t})$. The reconstructed text $\hat{T}$ is generated by decoding the fused feature $f_{f}$ using a BERT-based [11] text decoder $\hat{T}=D_{t}(f_{f})$. The MLM loss for each masked word $t\in\left[T_{m},T_{a}\right]$ is defined as $\mathcal{L}^{t}_{\mathrm{MLM}}(\hat{t},t)=-\mathrm{log}(P(\hat{t}=t))$.
@@
Moreover, we simply categorize $T_{a}$ into two groups: negative descriptions $T_{neg}$ and others $T_{oth}$, with $N^{T}_{neg}$ and $N^{T}_{oth}$ words, respectively. Please note that since the expressions of positive descriptors are diverse, we divide them into others $T_{oth}$. To better facilitate the pre-training model to learn precise disease semantics, we focus on modeling the minority-positive disease descriptions instead of majority-negative words, technically, we multiply the loss of $T_{neg}$ by a small number $\lambda_{neg}$, and the loss of $N^{T}_{oth}$ by $\lambda_{oth}$, where
@@
According to the statistical results in Fig. 2 (b)., the most negative descriptor “there is no” occurs approximately 20×20$\times$20 × more often than the most positive descriptor “there may be.”, so we set $\lambda_{neg}=0.05$. Please note that $\lambda_{oth}$ is important to make sure that the model pays an enough attention to accurately reconstruct the critical “descriptive words” $T_{a}$. The final loss of entity-centered context-aware MLM is defined:
@@
Method ChestX-ray14 (AUC) CheXpert (AUC) RSNA (AUC) COVIDx (ACC)